### v1

# Latent Neural Operator for Solving Forward and Inverse PDE Problems

Abstract
Neural operators effectively solve PDE problems from data without knowing the explicit equations, which learn the map from the input sequences of observed samples to the predicted values. Most existed works build the model in the original geometric space, leading to high computational costs when the number of sample points is large. We present the Latent Neural Operator (LNO) solving PDEs in the latent space. In particular, we first propose Physics-Cross-Attention (PhCA) transforming representation from the geometric space to the latent space, then learn the operator in the latent space, and finally recover the real-world geometric space via the inverse PhCA map. Our model retains flexibility that can decode values in any position not limited to locations defined in training set, and therefore can naturally perform interpolation and extrapolation tasks particularly useful for inverse problems. Moreover, the proposed LNO improves in both prediction accuracy and computational efficiency. Experiments show that LNO reduces the GPU memory by 50%, speeds up training 1.8 times, and reaches state-of-the-art accuracy on four out of six benchmarks for forward problems and a benchmark for inverse problem.
1 Introduction
Neural network approaches for partial differential equations (PDEs) attract increasing attentions in diverse scientific fields, for instance, meteorological prediction[1], industrial design[2], geological sensing[3] and environmental monitoring[4] to name a few. Compared with traditional numerical methods, e.g., finite element method[5], which demands large amount of computational resources and requires meticulous discretization involving a lot of specialized knowledge, neural networks provide a new paradigm for solving PDE-driven problems with better trade-off between accuracy and flexibility.
Neural PDE solvers can be divided into two main categories, physics-informed and operator-learning models. The physics-informed methods[6, 7, 8, 9, 10, 11] enforce the inductive bias of explicit PDE instances into loss function or model architecture, usually yielding solutions with high accuracy, but are less flexible as they require fresh training to find the solution for each new instance. In contrast, the operator learning methods[12, 13, 14, 15] adopt a data-driven paradigm to learn the implicit PDE constraint from pair of samples representing input-to-output functional mappings, providing better flexibility that can generalize to different initial/boundary conditions beyond the scope of the training data, but have yet large room to improve in the prediction accuracy.
Recently, Transformers[16] are prevailing in neural operator structures. Theoretically, attention mechanism was proved to be a special form of the integral kernel[15, 17] in neural operator, which naturally conforms to sequence-to-sequence characterization. Moreover, the fully-connected attention structure enables the model characterizes long-distance and quadratic relationship among sample points, which yields more accurate prediction than the vanilla MLP structure[13] but at an expense of increasing the computational complexity drastically, as the complexity is quadratic respect to the length of input sequence.
To alleviate the complexity, existing works employed linear attention mechanism [18, 19, 20, 21] speeding up the computation but sacrificing precision due to their limited expressive power. Another lines of works attempted to solve the PDE in latent space with small number of samples [22, 23, 24], which get rid of abundant sample points in the original geometric space and capture physical correlations in the tighter latent space.
In this work, we present the Latent Neural Operator (LNO) with particular effort on designing the Physics-Cross-Attension (PhCA) module to learn the latent space to optimize the trade-off between accuracy and flexibility, and reduce the computational complexity as well. Specifically, LNO first encodes the input sample sequence into a learnable latent space by PhCA. Then, the model learns the PDE operator in the latent space via a stack of Transformer layers. Finally, the predicted output function is decoded by the inverse PhCA, given any query locations. Unlike previous works that either pre-define the latent space as frequency domain [14], or switch back-and-forth between latent space and geometric space at each layer [23], the latent space of the LNO is learnable from data and the data are only transformed at the first and last layer.
The proposed PhCA module decouples the locations of input samples fed in the encoder and output samples fed in the decoder, which can predict values in any position not limited by the observation. Therefore, it can naturally perform interpolation and extrapolation tasks, which is fundamental for solving inverse problems. The proposed LNO improves in both prediction accuracy and computational efficiency. Experiments shows that LNO reduces the GPU memory by 50% and speeds up training 1.8 times. Moreover, it reaches state-of-the-art accuracy on four (Pipe, Elasticity, Darcy, Navier-Stokes) out of six benchmarks for forward problems and a benchmark for inverse problem on Burgers’ equation.
Finally, the main contributions of our LNO framework are summarized as follow.
  • Flexibility: We propose Physics-Cross-Attention (PhCA) module which decouples the locations of input and output samples and learns the latent space from data. We also build the Latent Neural Operator (LNO) model for solving both forward and inverse PDE problems.
  • Efficiency: The LNO model transforms the data to the latent space only once. Compared with other approaches that switch the latent space and geometric space in each Transformer layer, the proposed model reduces complexity drastically in term of GPU memory usage, training time and number of model parameters.
  • Accuracy: We obtain state-of-the-art results on the forward problems of Pipe, Elasticity, Darcy, Navier-Stokes, and the inverse problem of Burgers’ equation.
2 Related works
We first review two major classes of neural PDE approaches, physics-informed and operator-learning methods. Then, we briefly discuss the works using neural networks for solving inverse PDE problems.
2.1 Physics-Informed Machine Learning
Physics-informed machine learning methods incorporate prior knowledge of PDEs into loss function or model architecture. The Physics-Informed Neural Network (PINN)[7] and its variants are among the most classic methods. PINN integrates the differential equation, initial conditions and boundary conditions into the loss function, guiding the neural network to approximate the solution by ensuring its derivatives satisfies the equation and the output aligns with the initial and boundary conditions. Other methods attempt to encode physical priors into the model architecture, e.g., the Finite Element Neural Network (FINN)[8] and Physics-encoded Recurrent Convolutional Neural Network (PeRCNN)[10]. These methods achieve high precision, but require solving optimization to train the networks for each instance. In addition, in the case, where concrete PDE is unknown or not exact, it is challenging to adopt those methods, limiting the generalization ability and flexibility in practical data-rich situations.
2.2 Operator Learning
Operator learning methods aim to find a functional map from coefficients, forcing terms and initial conditions to the solution function by learning an approximating operator. These methods train neural networks to capture the correspondence between input and output functions from data without needing to know the underlying PDE explicitly. DeepONet[13, 25] first introduces approximating operators using neural networks.
A series of kernel integral operator forms were proposed to model the mappings between functions as neural operators[26, 14, 15]. Specifically, Galerkin Transformer[20] utilizes Galerkin-type attention mechanism as kernel integral operators. OFormer[27] improves Galerkin-type attention by decoupling observation positions and query positions through cross attention. GNOT[21] introduces heterogeneous cross attention to modulate various input conditions for query positions, and alleviates computational pressure of the large grid by linear-time-complexity attention mechanism. FactFormer[28] proposes axial factorized kernel integral, decomposing input functions into multiple one-dimensional sub-functions, effectively reducing computation. ONO[29] enhances the generalization of neural operators by adding orthogonality regularization in attention mechanisms.
Researchers also explored the idea of modeling in latent space. Transolver[23] employs Physics-Attention in each layer to map back-and-forth between geometric features and physical features. LSM[22] encodes the input function into the latent space using cross-attention, constructs orthogonal basis, and then decodes the solution back to the geometric space through cross-attention again. UPT[24] introduces encoding and decoding losses, compresses geometric point information into super-node information in latent space via graph neural network[30].
Compared with existing methods, we autonomously learn mappings between functions in real-world space and the latent space in an end-to-end fashion with the cross-attention mechanism, without manually constructed orthogonal basis nor additionally introduced loss.
2.3 Inverse Problem
Inverse problems of PDEs have extensive applications in fields involving sensing and reconstruction such as medical imaging and geological sensing. For instance, Transformer-based deep direct sampling methods[31] have been proposed for electrical impedance tomography, and convolutional network methods[32, 33, 34] are widely used in full waveform inversion.
Theoretical works on solving inverse problems have also been proposed. The reconstruction of diffusion fields with localized sources have been studied[35] from the perspective of sampling matrix in spatiotemporal domain. PINN[36, 37, 38] approximates the solution based on partially observed data to compute the unknown data. NIO[39] combines DeepONet[13] and FNO[14] to construct a framework for solving inverse problems. FINN[8, 40] sets boundary conditions as learnable parameters, and infer boundary values through backpropagation. We aim to unify the solution of forward and inverse problems in the latent space through the PhCA module, which decouples observation and prediction positions.
3 Method
We formally define both the forward and the inverse problems of PDEs. Then, we introduce our Latent Neural Operator (LNO) model. Finally, we discuss the core module that learns transformation between real-world space and latent space via cross-attention mechanism.
Figure 1: The overall architecture of Latent Neural Operator.
3.1 Problem Setup
We consider a partial differential equation (PDE) with boundary and/or initial conditions defined on $D\subseteq\mathbb{R}^{d}$ or $\mathbb{R}^{d}\times[0,T]$
$\mathcal{L}_{a}\circ u$ $=0,$ (1)
$u({x})$ $=b({x}),\quad{x}\in\partial D$
where $\mathcal{L}_{a}$ is an operator containing partial differentials parameterized by a set of coefficients $a$, and $u(x)$, $x\in D$ and $b({x})$ represent the solution to the PDE and the boundary/initial conditions respectively.
A conventional forward problem aims to find the solution $u(x)$, given the differential operator $\mathcal{L}_{a}$ in (1) with the explicit parameters $a$ and the boundary/initial conditions $b(x)$. For an inverse problem, given an observed set of partial data $\{(x,u(x))|\;x\in\tilde{D}\subset D\}$, the task is to infer the PDE that generates the observed data along with possible boundary/initial conditions or the parameters $a$ in the differential operator $\mathcal{L}_{a}$.
Both forward and inverse problems can be unified as an operator learning task fitting into a sequence-to-sequence translation framework. An operator model estimates the mapping $\mathcal{F}:h\mapsto g$ from an input function $h$ to an output function $g$ by the data-driven approach using paired training data of input-output sequences ($(\{h(y_{i})\}_{i},\{g(z_{j})\}_{j})$.
Specifically, for the forward problem, the inputs are samples of either boundary/initial function $b$ or PDE coefficient function $a$, and the output are the physical quantity $u$ evaluated at a grid of points. Conversely, for the inverse problem, the input are a few partially observation samples $\{(pos_{in}^{(i)},val_{in}^{(i)})\}^{N_{in}}=\{(x_{i},u(x_{i}))\}_{i=1}^{N_{in}}$, and the outputs are boundary/initial function (e.g., infer initial state) or the PDE coefficient function (also coined as system identification problem).
The challenge in solving forward and inverse problems of PDEs lies in the large lengths of input and output sequences. In terms of operator approximating, complex kernel integrals lead to high computational cost while simple kernel integrals lack in accuracy. In this work, we build accurate and computationally efficient model for both forward and inverse problem by learning the operator in the learnable latent space.
3.2 Latent Neural Operator
Overview
The proposed model of Latent Neural Operator (LNO) is applicable for both forward and inverse PDE problems. It consists of four modules: an embedding layer to lift the dimension of the input data, an encoder to transform the input data to a learnable latent space, a sequence of self-attention layers for modeling operator in the latent space, and a decoder to recover the latent representation to the real-world space. The overall architecture of LNO is shown in Figure 1
Embedding inputs
First, we embed the positions and physics quantities of the input sequence to a higher dimensional representation. It is implemented by two MLPs, where the first only embeds the position information and the second embeds both the location and physics quantities
${x}^{(i)}$ $={\rm trunk\text{-}projector}\big{(}{pos}_{in}^{(i)}\big{)},$ ${x}^{(i)}\in\mathbb{R}^{d\times 1}$
${y}^{(i)}$ $={\rm branch\text{-}projector}\big{(}{\rm concat}({pos}_{in}^{(i)},{val}_{in}^{(i)})\big{)},\quad$ ${y}^{(i)}\in\mathbb{R}^{d\times 1}.$
Analogous to DeepONet [13], we named these MLPs as trunk-projector and branch-projector respectively. The embedding operation helps to map the locations and physics quantities to the space where their relationships are easier to be captured.
Encoding to latent space
We seek the representation of the input function in a latent space, where the number $M$ of sample points in the latent space can be much smaller than the number $N_{in}$ of samples of raw inputs. We use the cross-attention to model space transformation, where the query $\boldsymbol{H}$, key $\boldsymbol{X}=[x^{(1)},x^{(2)},\ldots,x^{(N_{in})}]^{T}$ and value $\boldsymbol{Y}=[y^{(1)},y^{(2)},\ldots,y^{(N_{in})}]^{T}$ inputs corresponding to locations in the target latent space, embedding of locations in the source real-world space, and embedding of locations and physics quantities in the source real-world respectively
$\boldsymbol{Z}^{0}={\rm Physics\text{-}Cross\text{-}Attention}(\boldsymbol{H},\boldsymbol{X},\boldsymbol{Y})$ (2)
In contrast to FNO[14], where the latent space is pre-defined as the frequency domain, we learn the latent space from data. Therefore, the sample locations $\boldsymbol{H}$ as query in the cross-attention are also learnable parameters.
After encoding the input function into the latent space, the length of the sequence to be processed is reduced from $N_{in}$ to $M$. With $M\ll N_{in}$, extracting the feature of the input function in the latent space will be much more efficient than in the real-world geometry space.
Learning operator in the latent space
We model the operator of the forward/inverse PDE problem in the latent space, using a stack of Transformer blocks with self-attention mechanism
$\boldsymbol{Z}^{l}={\rm MLP}({\rm LayerNorm}(\boldsymbol{\hat{Z}}^{l}))+\boldsymbol{\hat{Z}}^{l}\text{ \quad with }\boldsymbol{\hat{Z}}^{l}={\rm Attention}({\rm LayerNorm}(\boldsymbol{Z}^{l-1}))+\boldsymbol{Z}^{l-1}$
where $l\in\{1,2,...,L\}$ is the index of Transformer block with $L$ being the number of the blocks. We experiment several implementations of attention mechanisms in the ablation study, and find the classical scaled dot-product attention [16] performs consistently well on various tasks.
Decoding to real-world space
Finally, we decode the output sequence from the latent sequence through Physics-Cross-Attention according to the query ${pos}_{out}$ of the output sample locations, and use another MLP to map the embbedings to the values of output function
${p}^{(j)}$ $={\rm trunk\text{-}projector}({pos}_{out}^{(j)}),$ ${p}^{(j)}\in\mathbb{R}^{d\times 1}$ (3)
$\boldsymbol{U}$ $={\rm Physics\text{-}Cross\text{-}Attention}(\boldsymbol{P},\boldsymbol{H},\boldsymbol{Z}^{L})$
${{val}}_{out}^{(j)}$ $={\rm MLP}({u}^{(j)}),$ ${u}^{(j)}\in\mathbb{R}^{d\times 1}$
During the inferring phase, the trunk-projector can accept inputs from any position, enabling LNO generalize to unseen region where the positions are not presented in training phase.
3.3 Physics Cross Attention
The Physics-Cross-Attension (PhCA) is the core module in the aforementioned latent neural operator (LNO) model, acted as the encoder and decoder which transform the data representation back-and-forth between the real-world space and the latent space. Motivated by the work [21] that learns the operator mapping cross domains using the cross-attention, we use the same cross-attention to model the transformation to/from the latent space, but set the latent space learnable rather than pre-defined, e.g., FNO [14].
In particular, the rows of the query $\boldsymbol{H}$ and key $\boldsymbol{X}$ matrices in (2) represent the embedded positions of samples in the target and source space respectively. A row of the value matrix $\boldsymbol{Y}$ is the embedding of both position and physics quantity of a sample in the source space. We design the latent space to be learned from data. Thus, the positions $\boldsymbol{H}$ of samples in the latent space should be set as learnable parameters as well. Then, the cross-attention of the encoder in (2) is simplified as
$\boldsymbol{Z}^{0}={\rm softmax}\Big{(}\tfrac{1}{\sqrt{d}}\boldsymbol{H}\boldsymbol{W}_{q}\boldsymbol{W}_{k}^{T}\boldsymbol{X}^{T}\Big{)}\boldsymbol{Y}\boldsymbol{W}_{v}={\rm softmax}\big{(}\boldsymbol{W}_{1}\boldsymbol{X}^{T}\big{)}\boldsymbol{Y}\boldsymbol{W}_{v},$ (4)
where we merge the learnable matrices of query $\boldsymbol{H}$, query weight $\boldsymbol{W}_{q}$ and key weight $\boldsymbol{W}_{k}$ as a single matrix $\boldsymbol{W}_{1}$.
Correspondingly, the cross-attention in the decoder has a learnable key matrix that represents sample positions in the latent space. Therefore, the decoder in  (3) is written as
$\boldsymbol{U}={\rm softmax}\Big{(}\tfrac{1}{\sqrt{d}}\boldsymbol{P}\boldsymbol{W}_{q}^{{}^{\prime}}\boldsymbol{W}_{k}^{{}^{\prime}T}\boldsymbol{H}^{T}\Big{)}\boldsymbol{Z}^{L}\boldsymbol{W}_{v}^{{}^{\prime}}={\rm softmax}(\boldsymbol{P}\boldsymbol{W}_{2}^{T})\boldsymbol{Z}^{L}\boldsymbol{W}_{v}^{{}^{\prime}}.$ (5)
In addition, utilizing the relationship of the encoder and decoder as mutually inverse transformations, we set $\boldsymbol{W}_{1}=\boldsymbol{W}_{2}$ to reduce the number of parameters, which also improves model performance experimentally. Practically, we generalize the linear projection $\boldsymbol{W}_{1},\boldsymbol{W}_{2}$ with an MLP for further improvement, which is named attention-projector as shown in Figure 2.
In the special case where the input only contains sample positions without physics quantities, our PhCA module is similar to the transformation of physics-aware token in Transolver [23] (only differing in normalizations). In general cases when the inputs inlcude both positions and physics quantities, the two approaches are different. Specifically, in Transolver, the key $\boldsymbol{X}$ in (4) and query $\boldsymbol{P}$ in (5) contain both positions and physics quantities information, whereas in PhCA, the key matrix $\boldsymbol{X}$ in the encoder and the query matrix $\boldsymbol{P}$ in the decoder only depend on the positions of the input and output respectively. The physics quantities are used only to produce the value matrix. Such decoupling property of values and locations in query and key enables LNO to predict values at different positions from input positions.
Figure 2: The working mechanism of Physics-Cross-Attention in encoder and decoder respectively.
4 Experiments
We first evaluate the accuracy for the forward problems on the six popular public benchmarks and for the inverse problem using Burger’s equation. Then, we measure the efficiency in term of the number of model parameters, GPU memory usage and the training time. Finally, we establish ablation study on the choices of different types of attention mechanisms and sample sizes in latent space. Additional experiments on scaling, sampling rate, and sharing weights between encoder and decoder are included in the appendix.
4.1 Accuracy for the forward problems
We conduct experiments for the forward problem on six benchmarks including Darcy and NS2d on regular grids (refer to section 5.2, 5.3 in [14] for details), and Airfoil, Elasticity, Plasticity and Pipe problems on irregular grids (refer to section 4.1–4.5 in [41] for details).
Table 1 shows a comprehensive comparison with previous works who reported their performance on the same benchmarks. Our LNO achieves best accuracy on four benchmarks: Darcy, NS2d, Elasticity and Pipe. On the Airfoil and Plasticity benchmarks, our LNO achieves performance close to the SOTA with only half computational cost discussed in Section 4.3. These results demonstrate the effectiveness of approximating the kernel integral operator in the latent space through PhCA module.
Table 1: Prediction error on the six forward problems. The best result is in bold. We reproduce the Transolver by implementing codes independently (marked with $\ast$) besides the results claimed in the original paper[23]. The column labeled D.C. indicates whether the observation positions and prediction positions are decoupled. Standard deviations are computed based on 5 independent trials.
Model D.C. Relative L2(×$\times 10^{-2}$)
Darcy NS2d Airfoil Elasticity Plasticity Pipe
FNO[14] N 1.08 15.56 / / / /
Geo-FNO[41] N 1.08 15.56 1.38 2.29 0.74 0.67
U-FNO[42] N 1.83 22.31 2.69 2.39 0.39 0.56
F-FNO[43] N 0.77 23.22 0.78 2.63 0.47 0.70
LSM[22] N 0.65 15.35 0.59 2.18 0.25 0.50
Galerkin[20] N 0.84 14.01 1.18 2.40 1.20 0.98
OFormer[27] Y 1.24 17.05 1.83 1.83 0.17 1.68
GNOT[21] Y 1.05 13.80 0.76 0.86 3.36 0.47
FactFormer[28] N 1.09 12.14 0.71 / 3.12 0.60
ONO[29] Y 0.76 11.95 0.61 1.18 0.48 0.52
Transolver[23] N 0.57 9.00 0.49 0.64 0.12 0.33
[…]
LNO(Ours) Y 0.49 ($\pm$0.01) 8.45 ($\pm$0.22) 0.51 ($\pm$0.05) 0.52 ($\pm$0.03) 0.29 ($\pm$0.03) 0.26 ($\pm$0.03)
4.2 Accuracy for the inverse problem
To demonstrate the flexibility of LNO, we design an inverse problem. Given a set of partially observed solution $u(x)$, the aim is to recovery the complete $u(x)$ in a larger domain. Specifically, we test on 1D Burgers’ equation
$\tfrac{\partial}{\partial t}u(x,t)=0.01\tfrac{\partial^{2}}{\partial x^{2}}u(x,t)-u(x,t)\tfrac{\partial}{\partial x}u(x,t),\quad x\in[0,1],\;t\in[0,1]$
$u(x,0)\sim\text{GaussianProcess}(0,\exp\big{(}-\tfrac{2}{pl^{2}}\sin^{2}(\pi||x-x^{{}^{\prime}}||\big{)},\quad u(0,t)=u(1,t)$
The ground-truth data is generated on a $128\times 128$ grid with the periodic boundary condition. Initial conditions are sampled from the Gaussian process with the periodic length $p=1$, scaling factor $l=1$.
The objective of this inverse problem is to reconstruct the complete solution $u(x)$ in the whole spatiotemporal domain $(x,t)\in[0,1]\times[0,1]$ based on sparsely random-sampled or fixed observation in the sub-spatiotemporal domain $(x,t)\in[0,1]\times[0.25,0.75]$, as illustrated in Figure 6 in the appendix.
Instead of naive approach which directly learns the mapping from partially observed samples in the subdomain to the complete solution in the whole domain, we propose a two-stage strategy. First, we train the model as a completer to interpolate the solution in the subdomain. Then, we train the model as a propagator to extrapolate from the subdomain to the whole domain. Since, the observation and the prediction samples locate in different positions, only the models with decoupling property can be used as the completer and propagator.
We compare the performance of our LNO with DeepONet[13] and GNOT[21] in two stages respectively. The results in Table 2 and Table 3 indicate that i) in the first stage, LNO performs significantly better than DeepONet and GNOT at different random observation ratio; ii) in the second stage, LNO performs better when the complete solution in the subdomain approaches the ground truth, and the performance degrades as the reconstruction error of the complete solution in the subdomain increases. This is reasonable since we train the propagator based on the ground truth in the sub domain. When there is a significant discrepancy between the ground truth and the prediction of completer, the model faces challenging distribution shifts. Nevertheless, when LNO is used both as the completer and propagator, it achieves the most accurate results for solving this inverse problem.
We also investigate the impact of different temporal and spatial sampling intervals on the accuracy of LNO in the fixed grid observation situation. See appendix B for details.
Table 2: Relative MAE of different completers in the subdomain in the 1st stage of the inverse problem with different settings of observation ratio.
Completer |||| Observation ratio 20% 10% 5% 1% 0.5%
DeepONet[13] 2.51% 2.59% 2.82% 3.25% 4.82%
GNOT[21] 1.12% 1.39% 1.62% 1.63% 2.56%
LNO(Ours) 0.60% 0.74% 0.77% 1.18% 2.05%
Table 3: The reconstruction error of different propagators in the 2nd stage of the inverse problem. Propagators trained to reconstruct the solution in the whole domain based on the ground truth (G.T.) of the subdomain and are tested using the output of different completers. Relative MAE of $t=0$ and $t=1$ is reported.
Propagator |||| Completer G.T. LNO GNOT DeepONet
10% 1% 10% 1% 10% 1%
DeepONet[13] 7.34% 8.01% 9.38% 9.09% 10.80% 11.14% 13.87%
GNOT[21] 5.45% 6.50% 8.07% 8.04% 9.91% 10.41% 13.45%
LNO(Ours) 3.73% 5.69% 7.72% 9.03% 10.98% 13.11% 15.50%
Table 4: Comparison of efficiency between LNO and Transolver on the six forward problem benchmarks. Three metrics are measured: number of parameters, cost of GPU memory and cost of training time per epoch.
Metric Model Darcy NS2d Airfoil Elasticity Plasticity Pipe
Paras Count(M) Transolver 1.91 5.33 1.91 1.91 1.91 1.91
LNO 0.76 5.08 1.36 1.42 1.36 1.36
Memory(GB) Transolver 17.11 17.17 4.49 1.48 18.41 5.94
LNO 5.75 7.58 2.47 1.39 7.16 2.89
Time(s/epoch) Transolver 88.68 107.62 19.49 5.66 83.43 25.56
LNO 38.98 57.83 9.35 5.33 41.62 14.13
4.3 Efficiency
We demonstrate improvement on efficiency of our proposed LNO compared to Transolver, which was considered the most efficient Transformer-based PDE solvers[23]. Results in Table 4 indicate that our method has fewer parameters, lower memory consumption, and faster training speed. Roughly speaking, LNO reduces the parameter count by an average of 30% and memory consumption by 50% compared to Transolver across all benchmarks, also with a 1.8x training speed up.
The improvement of LNO stems from the idea of solving PDEs in the latent space. Given $N$ input points in real-world space, if both of Transolver and LNO are consisted of $L$ Transformer blocks, the Transolver, which applies Physics-Attention in each block, has the time complexity of $O(LMN+LM^{2})$, where $M$ is the number of slices. Our LNO, which applies Physics-Cross-Attention only in the encoder and decoder, has the time complexity of $O(MN+LM^{2})$, where $M$ is the sample size in latent space.
4.4 Ablation Study
Necessity of decoupling
Decoupling the values and locations as well as the observation positions and prediction positions not only guarantees the flexibility but also improves the accuracy. To study the effect of decoupling property, we replaces the PhCA in LNO by Physics-Attention proposed in Transolver[23]. In this way, the query and keys are calculated from the same embedding of both ${pos}_{in}$ and ${val}_{in}$ of input sequence. The ${pos}_{out}$ which not presents in the input sequence can not generate queries and keys. So the observation positions and prediction positions must be kept the same. We compare the performance of LNO with PhCA or Physics-Attention on the forward problem benchmarks, and the results in the first and last rows of Table 5 verify the necessity of decoupling.
Expressiveness of quadratic attention
We study the impact on accuracy using different attention mechanisms in the latent space. We replace the scaled dot-product attention in LNO with Galerkin-type attention used in Galerkin Transformer[20], efficient attention[18] used in GNOT[21], and Nystrom attention[19] used in ONO[29]. The comparison results in Table 5 show that quadratic attention such as scaled dot-product attention[16] is more expressive than linear attention, which helps improve the accuracy. Many previous methods have employed different linear attentions[20, 21, 29] to reduce the computational costs at the expense to reduce the accuracy. Since the length of the latent sample sequence is much smaller than the inputs and outputs in real-world space, therefore, applying quadratic attention in latent space does not incur significant overhead.
Table 5: The ablation results using different attention implementation. P.A. stands for Physics-Attention. E.A. stands for Efficient Attention, N.A. stands for Nystrom Attention and G.A. stands for Galerkin-type Attention.
Model Relative L2(×$\times 10^{-2}$)
Darcy NS2d Airfoil Elasticity Plasticity Pipe
LNO with P.A.[23] 0.53 22.65 0.74 0.88 0.49 0.40
LNO with E.A.[18, 21] 0.63 20.18 0.59 0.54 0.44 0.36
LNO with N.A.[19, 29] 0.67 23.03 0.63 0.79 0.30 0.48
LNO with G.A.[20] 0.71 22.44 0.89 0.93 0.43 0.39
LNO 0.49 8.45 0.51 0.52 0.29 0.26
Sample size in latent space
We investigate the impact of sample size in latent space on LNO for solving both the forward and inverse problems. Results are shown in Figure 3. In the forward problems, as the sample size $M$ in latent space increases, the performance of LNO gradually improves to a saturation level, followed by a gentle decline on the benchmarks of Airfoil, Elasticity, Plasticity, and Pipe. However, on benchmarks of Darcy and NS2d, the performance of LNO continues to improve till $M=512$. This may be attributed to the former set of benchmarks containing only location information, while the latter set involves both location and physics quantity information, which requires more samples to represent. In the inverse problem, the performance of completer and propagator are similar to the trends in the forward problems.
Figure 3: Sample size in latent space v.s. accuracy. (a) Forward problems on Airfoil, Elasticity, Plasticity, Pipe. (b) Forward problems on Darcy, NS2d. (c) Inverse Problem.
5 Conclusions
The paper explores the feasibility of characterizing PDEs in a learnable latent space. We present Physics-Cross-Attension (PhCA) for encoding inputs into the latent space and for decoding outputs. We build the Latent Neural Operator (LNO) based on PhCA and validate its flexibility, efficiency and accuracy for solving both forward and inverse problems. In this work, no prior is imposed on the latent space, but, in future, it is worthy to explore whether physical priors, e.g. symmetry or disentanglement, can help gain further improvements, and explore generalization ability of PhCA crossing domains and re-usability for multiple types of PDEs.
Appendix A Scaling Experiments
We conduct scaling experiments to investigate the impact of model depth (number of stacked Transformer blocks) and model width (dimension of tokens in Transformer blocks) on the performance of LNO.
As shown in Figure 4, with model depth increasing, the performance of LNO initially improves and then declines, reaching its optimal performance at 4 or 8. This suggests that deeper LNO encounters optimization challenges in solving PDEs in the latent space. Simply stacking more Transformer blocks does not necessarily lead to better accuracy.
As shown in Figure 5, although the performance of LNO continues benefiting from increasing model width, it gradually exhibits a saturation trend when the dimension of tokens reaches as large as 256.
Figure 4: The influence of model depth on the performance of LNO across six forward problem benchmarks.
Figure 5: The influence of model width on the performance of LNO across six forward problem benchmarks.
Appendix B Observation Ratio and Locations for the Inverse PDE Task
We investigate the impact of different temporal and spatial sampling intervals on the accuracy of LNO when solving inverse problem in the fixed observation situation. As shown by Figure 8, it can be found that i) the inverse problem can still be solved accurately even when the temporal and spatial sampling intervals are both as large as 16 (approximately only 0.4% observation ratio); ii) the increase of the spatial sampling interval significantly compromises the solution accuracy compared to that of the temporal sampling interval; iii) spatial undersampling can hardly be compensated for through temporal oversampling.
Figure 6: Visualization of Burgers’ equation with different observation situations in different regions. We propose a two-stage strategy. First, we interpolate the solution in the subdomain. Then, we extrapolate from the subdomain to the whole domain.
Figure 7: Visualization of Burgers’ equation in the fixed observation situation with different temporal and spatial sampling intervals.
Figure 8: Accuracy of solving inverse problem at different temporal and spatial sampling intervals. (a) Accuracy of LNO as completer. (b) Accuracy of LNO as propagator based on the results of corresponding completer.
Appendix C Additional Ablation Study
We conduct additional ablation study to explore the performance of LNO when ${W}_{1}\neq{W}_{2}$ in (4) and (5). In this way, we replace the linear projection $W_{1}$ and $W_{2}$ with two individual MLPs. Results of comparison between LNO using shared MLPs and LNO using non-shared MLPs in PhCA encoder and decoder are shown in Table 6. It can be found that LNO performs better on most benchmarks (except Plasticity) under the shared situation.
Table 6: The ablation results on six forward problem benchmarks. Relative L2 is recorded. Smaller values represent better performance. The better result is in bold.
Model Relative L2(×$\times 10^{-2}$)
Darcy NS2d Airfoil Elasticity Plasticity Pipe
LNO(non-shared) 0.54 10.67 0.53 1.09 0.26 0.36
LNO 0.49 8.45 0.51 0.52 0.29 0.26
Appendix D Full Implementation Details
The full implementation details of LNO in both forward and inverse problems are shown in Table 7. All the experiments are conducted on single or two NVIDIA RTX 3090 GPUs.
For the forward problems, we train the models using relative L2 error for 500 epochs. AdamW optimizer[44] and OneCycleLr scheduler[45] are applied with initial learning rate at $10^{-3}$. We reproduced the results of Transolver[23] following implementations in appendix B.3 in [23] and also excerpt the results of other methods from Table 2 in [23]. The data splits are the same as in appendix B.1 in [23].
For the inverse problem, we train the models using mean squared error and StepLr scheduler as well. For DeepONet[13], we use 4 Transformer blocks with 128-dimensional tokens as the branch net and 3 fully-connected layers as the trunk net. For GNOT[21], we set the number of Transformer blocks as 4, dimension of tokens as 96, and number of experts as 1. We use 4096 samples for training, 128 samples for validating and 128 samples for testing.
Table 7: The hyperparameters and training configuration of LNO in solving the forward and inverse problems in Section 4.1 and 4.2
Benchmark Layers Dim Size Heads Batch Epoch Loss Optimizer Scheduler
Darcy 4 128 256 8 4 500 rL2 AdamW[44] OneCycleLR[45]
NS2d 8 256
Airfoil 8 128
Elasticity 8 128
Plasticity 8 128
Pipe 8 128
Completer 4 96 256 8 4 500 MSE AdamW StepLR
Propagator
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We have clearly claimed the main contributions and scope of our work in the abstract and introduction, which can be reflected by the demonstration of methods in Section 3.2, 3.3 and experimental results in Section 4.1, 4.2, 4.3, 4.4.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We have discussed the limitations of our proposed LNO in Appendix, and potential approach to further improve the accuracy in conclusion seciton. A. The scalability of the model is lacking. Simply increasing the number of Transformer blocks does not always improve the model performance. In the conclusion section, we pose some future study on imposing additional prior on the latent space to expect further improvement.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Our work does not include theoretical result.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide detailed descriptions of our methods in Section 3.2 and Section 3.3, and full implementation details regarding the experiments in Appendix D. Additionally, we also provide our code and running instructions in the supplementary material.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have provided our code along with detailed instructions on the runtime environment, data access, data processing, and experiment conducting in the supplementary material documentation.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have provided the full implementation details in Appendix D including data splits, hyperparameters, optimizer, scheduler, epoch, etc. Also, the ablation studies in Section 4.4 and scaling experiments in Appendix A have shown the reasonableness of our choice for the hyperparameters.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We repeat all the forward problem experiments five times, and calculate the standard deviations of our proposed LNO, as shown in Table 1.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have provided a comprehensive summary of computational resources required for all the experiments in Appendix D, along with detailed records of memory, execution time, etc., in Section 4.3.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS Code of Ethics carefully and made sure our research conform with it in every respect.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our work is foundational research about solving forward and inverse problems of PDEs, which is not tied to particular applications or deployments. And there is no direct path to any negative applications.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our work primarily focuses on solving the forward and inverse problems of PDEs, which does not pose such risks of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have utilized publicly available datasets provided in [14] and [41], as mentioned in Section 4.1. The URLs of the assets are also provided in the supplementary material.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide the code of our proposed LNO in the form of an anonymous zip file. Detailed documentation is also included to explain how to use our code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.